# WISE: Web Information Satire and Fakeness Evaluation

### v1

@@ 3.3.5. Statistical Testing
We perform rigorous statistical testing to compare model performance across all evaluated models: (1) paired t-tests on 5-fold Macro-F1 scores with Benjamini-Hochberg correction for multiple comparisons to control false discovery rate, (2) McNemar tests on aggregated 2×2 contingency tables per fold to assess systematic differences in error patterns, (3) DeLong tests for ROC-AUC differences to compare discriminative capabilities, and (4) 95% confidence intervals for all metrics using t-distribution. These tests provide statistical rigor to our comparisons and identify significant performance differences between models. The statistical testing framework ensures that reported performance differences are not due to random variation but reflect genuine model capabilities.
@@ 3.3.6. Reproducibility
We use random seed 42 for data splits, model initialization, and training to ensure reproducibility. All experimental configurations are specified in YAML configuration files, with the default configuration used for all experiments. Experiments were conducted on a MacBook Pro with Apple M4 Pro chip and 24 GB unified memory, using PyTorch with MPS (Metal Performance Shaders) acceleration. All experimental configurations, source code, and analysis results are maintained in a version-controlled repository [^11] Repository link masked for review. under the MIT license, with full environment specifications (Python version, library versions, hardware details) logged in the run manifest. Model checkpoints, tokenizers, and training logs are saved for each fold, enabling full reproducibility and post-hoc analysis.
4. Results and Analysis
@@
Table 5 presents per-class performance metrics, revealing how each model handles satire and fake news classification. RoBERTa-base shows higher recall for satire (92.17%) but lower precision (84.40%), while BERT-base demonstrates more balanced performance across both classes. MiniLM shows the most balanced per-class performance, with F1-scores of 87.49% for satire and 87.66% for fake news, indicating effective handling of both classes without significant bias.
@@ 4.5. Model Architecture Insights and Deployment Recommendations
The performance variations can be attributed to fundamental architectural differences among the evaluated models. MiniLM’s superior accuracy (87.58%) stems from its sophisticated knowledge distillation approach, which effectively preserves the linguistic understanding capabilities of larger transformer models while maintaining computational efficiency. RoBERTa-base’s highest ROC-AUC (95.42%) reflects its optimized training procedures, including dynamic masking and improved optimization, which enhance the model’s ability to capture nuanced linguistic patterns. BERT-base’s strong performance validates the effectiveness of the original transformer architecture for this task. DistilBERT’s competitive performance demonstrates that knowledge distillation can effectively preserve model capabilities while significantly reducing computational requirements.
The performance differences between models are statistically meaningful, with MiniLM achieving the highest accuracy among all models. The performance gap between the best baseline (RoBERTa-base) and the best lightweight model (MiniLM) is minimal (0.22 percentage points in favor of MiniLM), indicating that lightweight models can match or exceed baseline performance for this task. Based on the comprehensive evaluation in Tables 4 and 5, we identify distinct deployment scenarios with corresponding model recommendations:
  • Maximum Accuracy: For scenarios requiring the highest classification accuracy, MiniLM represents the optimal choice, achieving 87.58% accuracy and 94.52% ROC-AUC while being more efficient than baselines.
  • Maximum Discriminative Power: For scenarios requiring the highest ROC-AUC, RoBERTa-base is optimal, achieving 95.42% ROC-AUC and 87.36% accuracy.
  • Balanced Performance: BERT-base offers strong performance (87.17% accuracy, 94.88% ROC-AUC) with well-established architecture and extensive community support.
  • Efficiency-Accuracy Trade-off: DistilBERT provides an excellent balance, achieving 86.28% accuracy and 93.90% ROC-AUC while being significantly smaller and faster than baseline models, making it ideal for resource-constrained deployments.
@@ 5.2. Model Architecture Insights
The differences in performance among the models can be attributed to their architectural design choices. MiniLM’s superior accuracy (87.58%) stems from its sophisticated knowledge distillation approach, which effectively preserves the linguistic understanding capabilities of larger transformer models while maintaining computational efficiency. This design enables MiniLM to achieve the highest accuracy among all models, demonstrating that knowledge distillation can create models that match or exceed baseline performance. RoBERTa-base’s optimized training procedures, including dynamic masking and improved optimization, enhance the model’s ability to capture nuanced linguistic patterns, enabling it to achieve the highest ROC-AUC (95.42%) and strong accuracy (87.36%). BERT-base’s strong performance (87.17% accuracy, 94.88% ROC-AUC) validates the effectiveness of the original transformer architecture for this task. DistilBERT’s knowledge-distillation approach proves effective in preserving model capabilities while significantly reducing computational requirements. DistilBERT achieves 86.28% accuracy and 93.90% ROC-AUC, demonstrating that knowledge distillation can effectively compress models without substantial performance degradation. The minimal performance gap (0.22 percentage points in favor of MiniLM) between baseline and lightweight models indicates that lightweight architectures can achieve competitive or superior performance for this task. In summary, our observations indicate that compact models can perform exceptionally well, with MiniLM achieving the highest accuracy. Maintaining sufficient representational capacity through knowledge distillation (as in MiniLM and DistilBERT) is important for capturing the differences between satirical and fake content. The success of MiniLM and DistilBERT suggests that knowledge distillation is an effective strategy for creating efficient models that retain or exceed the performance of larger architectures.

### v2

@@ 3.3.5. Statistical Testing
We perform rigorous statistical testing to compare model performance across all evaluated models: (1) paired t-tests on 5-fold Macro-F1 scores with Benjamini-Hochberg correction (Benjamini and Hochberg, 1995) for multiple comparisons to control false discovery rate, (2) McNemar tests on aggregated 2×2 contingency tables per fold to assess systematic differences in error patterns, (3) DeLong tests for ROC-AUC differences to compare discriminative capabilities, and (4) 95% confidence intervals for all metrics using t-distribution. These tests provide statistical rigor to our comparisons and identify significant performance differences between models. The statistical testing framework ensures that reported performance differences are not due to random variation but reflect genuine model capabilities.
@@ 3.3.6. Reproducibility
We use random seed 42 for data splits, model initialization, and training to ensure reproducibility. All experimental configurations are specified in YAML configuration files, with the default configuration used for all experiments. Experiments were conducted on a MacBook Pro with Apple M4 Pro chip and 24 GB unified memory, using PyTorch with MPS (Metal Performance Shaders) acceleration. All experimental configurations, source code, and analysis results are maintained in a version-controlled repository [^11] https://github.com/gauravfs-14/wise under the MIT license, with full environment specifications (Python version, library versions, hardware details) logged in the run manifest. Model checkpoints, tokenizers, and training logs are saved for each fold, enabling full reproducibility and post-hoc analysis.
4. Results
@@
Table 5 presents per-class performance metrics, revealing how each model handles satire and fake news classification. RoBERTa-base shows higher recall for satire (92.17%) but lower precision (84.40%), while BERT-base demonstrates more balanced performance across both classes. MiniLM shows the most balanced per-class performance, with F1-scores of 87.49% for satire and 87.66% for fake news, indicating effective handling of both classes without significant bias. Although performance differences across models are within  5%, this does not imply trivial classification. Error overlap analysis reveals that many samples misclassified by weaker models are also challenging for larger baselines, indicating intrinsic ambiguity rather than model capacity alone.
@@ 4.5. Model Architecture Insights and Deployment Recommendations
Performance differences in Tables 4 and 5 align with how much representational capacity each model retains under compression. MiniLM achieves the best overall accuracy (87.58%), marginally exceeding the strongest baseline RoBERTa-base (87.36%) by 0.22 percentage points, while RoBERTa-base provides the strongest ranking quality across thresholds with the highest ROC-AUC (95.42%). DistilBERT offers a strong accuracy–efficiency compromise, reaching 86.28% accuracy with a ROC-AUC of 93.90%. Smaller architectures (e.g., TinyBERT and Small-BERT-L2) exhibit a larger performance drop, consistent with their more limited capacity.
These results translate into practical deployment choices:
  • Maximum Accuracy: MiniLM (87.58% accuracy, 94.52% ROC-AUC) is the best choice when top-line accuracy is the primary objective.
  • Maximum Discriminative Power: RoBERTa-base (95.42% ROC-AUC, 87.36% accuracy) is preferred when threshold tuning and ranking quality matter most.
  • Efficiency-Accuracy Trade-off: DistilBERT (86.28% accuracy, 93.90% ROC-AUC) is a strong option for resource-constrained or latency-sensitive deployments.
  • Stable Baseline Reference: BERT-base (87.17% accuracy, 94.88% ROC-AUC) remains a reliable baseline with well-characterized behavior.
@@ 5.2. Model Architecture Insights
The benchmark suggests that distillation-based compression can preserve the cues required to separate satirical intent from deceptive framing, but only when the student model retains sufficient depth and attention capacity. MiniLM and DistilBERT exemplify this pattern: both are distilled models, and both remain close to the baselines, with MiniLM slightly leading in accuracy (87.58%) and DistilBERT maintaining a competitive profile (86.28%). In contrast, the smallest variants (e.g., 2-layer or very low hidden-size architectures) show a clearer degradation, indicating that aggressive compression likely removes higher-order signals useful for detecting irony, exaggeration, and incongruity.
RoBERTa-base achieves the highest ROC-AUC (95.42%), which indicates stronger separation between classes across decision thresholds even when its point accuracy is slightly below MiniLM. This behavior is consistent with training and pretraining choices that improve robustness and ranking quality, which is particularly relevant when systems must operate at different operating points (e.g., prioritizing recall of fake news vs. avoiding false alarms on satire). BERT-base remains a strong and stable reference, falling within the same top tier and providing a dependable comparison point.
Overall, the evidence supports two practical takeaways. First, distillation is a viable path to deployable misinformation classifiers without materially sacrificing performance, provided the compressed model preserves adequate representational capacity. Second, model selection should reflect the operating objective: MiniLM when top-line accuracy is the goal, RoBERTa-base when threshold flexibility and ranking quality are paramount, and DistilBERT when inference efficiency is the binding constraint.
@@ 5.3. Practical Implications
5.4. Limitations
Our study focuses on title-level binary classification on a single dataset, which may not capture nuanced misinformation in long-form articles or multimodal settings. Additionally, fixed hyperparameters were used for fair comparison, possibly preventing individual models from achieving optimal performance. Finally, results may not generalize across platforms beyond Reddit. Future work will explore cross-domain evaluation, classical baselines, and full-text multimodal modeling.